Two points charges are placed on the x-axis: a – 4- µC charge at x = 0 and a + 3- µC charge at x = 80 cm. At
what point(s) in the vicinity of these two charges can a + 5- µC charge be placed to that it experiences no
resultant force? ((5.176 m)
1
Expert's answer
2020-05-12T11:05:07-0400
"k\frac{q_1q_3}{x^2}+k\frac{q_2q_3}{(x-0.8)^2}=0"

"\frac{-4}{x^2}+\frac{3}{(x-0.8)^2}=0"

"x=5.971\ m"